OpenStudy (anonymous):

(Newton's Law of Motion) A car whose mass including the driver is 1600 kg has a maximum acceleration of 1.2 m/s^2. If three 80 kg passengers are also in the car, find its maximum acceleration then.

OpenStudy (anonymous):

F=ma =1600(1.2) =1920N with the same force: 1920=(1600+3x80) a a=1920/1780 =1.08ms^-2